# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v3

@@ Abstract
This paper presents UniVST, a unified framework for localized video style transfer based on diffusion model. It operates without the need for training, offering a distinct advantage over existing diffusion methods that transfer style across entire videos. The endeavors of this paper comprise: (1) A point-matching mask propagation strategy that leverages the feature maps from the DDIM inversion. This streamlines the model’s architecture by obviating the need for tracking models. (2) A training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels. This balances content fidelity and style richness, mitigating the loss of localized details commonly associated with direct video stylization. (3) A sliding-window consistent smoothing scheme that harnesses optical flow within the pixel representation and refines predicted noise to update the latent space. This significantly enhances temporal consistency and diminishes artifacts in stylized video. Our proposed UniVST has been validated to be superior to existing methods in quantitative and qualitative metrics. It adeptly addresses the challenges of preserving the primary object’s style while ensuring temporal consistency and detail preservation. Our code is available at https://github.com/QuanjianSong/UniVST.
@@
Diffusion Model, Training Free, Video Style Transfer.
@@ 1 Introduction
Video editing has greatly improved thanks to the use of diffusion methods [1, 2, 3]. T2V-Zero [4] changes self-attention mechanisms to cross-frame attention. Tune-A-Video customizes video editing by adjusting attention weights [5]. Fate-Zero [6] keeps the video content intact via information from the inversion process. Animate-Zero [7] and VideoBooth [8] add layers that stress time.
@@
Figure 1: Existing methods suffer from (a) lack of fine-grained control, (b) balance between content fidelity and style richness, and (c) temporal inconsistency. The last row shows the visualization of temporal consistency using MimicMotion’s Y-T slices [9] for both the existing method [10] and our proposed UniVST.
@@
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses image or text to direct the style. StyleCrafter [11] personalizes video style with extra training adapter; VideoBooth [8] uses image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from text descriptions. The second method includes all-encompassing video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, good for thematic expression but sometimes lacking in precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables aried scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
@@
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in a blurring and alter the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As seen in [10, 13], merely extending image stylization techniques to video can lead to inconsistencies across frames, manifesting as flickering and artifacts.
@@
In this paper, we propose UniVST, solving the above issues in a unified framework for localized [^1] Here, “localized” indicates a particular part of the video. video style transfer in a training-free manner. We inject the latent information from the content branch into the editing branch using mask, enabling the model to preserve the primary object’s style. Unlike in the image domain, where masks can be readily obtained using additional models such as those in [14, 15], the video domain presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. To address this, we introduce a point-matching mask propagation strategy leveraging feature maps from the DDIM inversion process [16] to capture correlations. Inspired by the image stylization model [17], replacing the key and value in the editing branch with those from style branch can facilitate style transfer. However, we found that directly applying this method in video style transfer can lead to the loss and blurring of local details. Based on this foundation, we develop a training-free AdaIN-guided style transfer mechanism, which functions at both the latent and attention levels. It adeptly balances content fidelity with style richness. Detailed comparison of results is presented in Fig. 4.
Our survey shows that existing methods often utilize optical flow information to achieve temporal consistency. For instance, Ground-A-Video [2] uses optical flow to smooth initial noise, ensuring its temporal coherence, while Flatten [1] integrates optical flow into the attention process to maintain frame consistency. Building on these insights, we present a sliding-window consistent smoothing scheme that employs optical flow within the pixel representation and refines predicted noise to update the latent space. As illustrated in Fig. 1, our method reduce flickering and artifacts while enhances the temporal consistency of the edited video. Overall, our major contributions in this paper are as follows:
@@
  • We introduce, to the best of our knowledge, the first localized video style transfer framework, featuring a novel point-matching mask propagation strategy.
@@
  • We develop a training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels. It effectively harmonizes content fidelity and style richness throughout the transfer process.
@@
  • We present a sliding-window consistent smoothing scheme based on optical flow, which adeptly upholds temporal consistency during the video transfer process.
@@
  • Extensive experiments demonstrate that our framework outperforms several state-of-the-art methods in both qualitative and quantitative metrics.
@@ 2 Related Work
Image Style Transfer. RSCT [18] calls for the separate foreground and background style transfer, and combines them to ensure a consistent style. DiffStyle [19] employs LoRA [20] to refine image prompts for stylization, while ArtBank [21] enhances text embeddings through fine-tuning image prompts. StyleID [17] leverages key-value replacement coupled with an initial latent AdaIN [22] to preserve content fidelity. Z* [23] rearranges attention to integrate content with style. InstanceStylePlus [24] utilizes adapters to safeguard content retention and enhance stylistic expression.
Video Style Transfer. OCD [25] integrates object-centric sampling and merging, expediting the editing process. StyleCrafter [11] refines adapters using image prompts based on the T2V model. Style-A-Video [26] enhances the stylistic alignment through strategic style guidance. Diffutoon [12] pioneers primary and editing branches for multi-segment editing in video colorization, employing FastBlend [27] to uphold temporal continuity. BIVDiff [10] combines frame-by-frame image editing with temporal consistency modeling. AnyV2V [13] follows a similar initial editing strategy, subsequently leveraging an I2V model [28] to propagate changes across the video sequence.
Figure 2: Overall framework. It is structured around three main components: (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Video Style Transfer(Attention-shift and latent-shift) and (3) Sliding-Window Consistent Smoothing.
@@ 3 Preliminary
Latent Diffusion Model. LDM [29] adds and removes noise in a low-dim space by using the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a Gaussian noise $Z_{T}$, DDIM denoising can be formulated as [3]:
@@
where ${Z}_{t\rightarrow 0}$ is an estimation of $Z_{0}$ at time step $t$:
@@
where $\alpha_{t}$ is a parameter and $C=\phi$ for editing integrity.
@@
where $A_{t}$ and $B_{t}$ are functions of time $t$. During editing, the initial noise $Z_{T}$ that aligns with the original video distribution $Z_{0}$ can be obtained using this equation.
3D UNet Extension. For video tasks, LDM requires a 3D UNet extension. Following [31, 1], the $3\times 3$ convolutional kernel in the convolutional blocks is expanded to $1\times 3\times 3$. To achieve frame parallelization, the feature shape $Z_{t}$ is transformed from $\mathbb{R}^{b\times c\times f\times h\times w}$ to $\mathbb{R}^{(bf)\times c\times h\times w}$ before self-attention and cross-attention operations. To facilitate frame interaction, we follow T2V-Zero [4] to modify the self-attention to cross-frame attention :
$\begin{array}[]{c}Q^{i}=W^{query}Z^{i},\quad K^{i}=W^{key}\left[Z^{{1}},Z^{{i-1}}\right],\\ V^{i}=W^{value}\left[Z^{{1}},Z^{{i-1}}\right],\end{array}$ (4)
where $Z^{i}$ denotes the feature of the $i$-th frame, and $W^{query}$, $W^{key}$, $W^{value}$ represent the respective mapping matrices, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicating concatenation.
Algorithm 1 Point-Matching Mask Propagation
1:Input: $M_{1}$ Mask for the first frame.
2:Input: $F_{t_{0}}$ Feature map from upsampling block-2.
@@
6:Output: $\{M_{i}\}_{i=1}^{N}$ Masks for all frames.
7:Initialize: previous_features [$\leftarrow$[\,]← [ ] Initialize empty list for previous features.
8:Initialize: previous_masks [$\leftarrow$[\,]← [ ] Initialize empty list for previous masks.
9:Initialize: first_feature [$\leftarrow$[\,]← [ ] Initialize empty list for first feature.
@@
11:    $fore\_index\leftarrow\text{where}(M_{i-1}=1)$
12:    $back\_index\leftarrow\text{where}(M_{i-1}=0)$
13:    $fore\_index\leftarrow\text{random\_sample}(fore\_index,r\cdot\frac{\lvert\text{fore}\rvert}{\lvert\text{fore}\rvert+\lvert\text{back}\rvert})$
14:    $back\_index\leftarrow\text{random\_sample}(back\_index,r\cdot\frac{\lvert\text{back}\rvert}{\lvert\text{fore}\rvert+\lvert\text{back}\rvert})$
15:    $current\_index\leftarrow\text{concat}(fore\_index,back\_index)$
16:    $current\_feature\leftarrow F_{t_{0}}^{i}[current\_index]$
@@
18:       first_feature.append($current\_feature$) Store first frame.
19:       $M_{i}\leftarrow M_{1}$
@@
22:          previous_features.pop(0) Remove the oldest frame feature.
23:          previous_masks.pop(0) Remove the oldest frame mask.
@@
26:       $M_{i}\leftarrow\text{KNN}(f,F_{t_{0}}^{i},\text{concat}(\text{previous\_masks},M_{1}),k)$
27:       previous_features.append($current\_feature$) Store current frame.
28:       previous_masks.append($M_{i}$) Store current mask.
@@
31:Return: $\{M_{i}\}_{i=1}^{N}$
4 Localized Video Style Transfer
Given an original video comprising $N$ frames $\{I^{i}\}_{i=1}^{N}$ and a reference style image $I^{s}$, our objective is to transform the original video into a new sequence $\{\mathcal{I}^{i}\}_{i=1}^{N}$. To achieve this, we propose UniVST, which integrates the style of the reference image into the original video while preserving the style of the primary objects within the video unchanged. As shown in Fig. 2, it mainly includes Point-Matching Mask Propagation, AdaIN-Guided Video Style Transfer, and Sliding-Window Consistent Smoothing three components.

### v4

@@ Abstract
This paper presents UniVST, a unified framework for localized video style transfer based on diffusion models. It operates without the need for training, offering a distinct advantage over existing diffusion methods that transfer style across entire videos. The endeavors of this paper comprise: (1) A point-matching mask propagation strategy that leverages the feature maps from the DDIM inversion. This streamlines the model’s architecture by obviating the need for tracking models. (2) A training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels. This balances content fidelity and style richness, mitigating the loss of localized details commonly associated with direct video stylization. (3) A sliding-window consistent smoothing scheme that harnesses optical flow within the pixel representation and refines predicted noise to update the latent space. This significantly enhances temporal consistency and diminishes artifacts in stylized video. Our proposed UniVST has been validated to be superior to existing methods in quantitative and qualitative metrics. It adeptly addresses the challenges of preserving the primary object’s style while ensuring temporal consistency and detail preservation. Our code is available at https://github.com/QuanjianSong/UniVST.
@@
Localized Video Style Transfer, Diffusion Models, Training-free.
@@ 1 Introduction
Video editing has greatly improved thanks to the use of diffusion models [1, 2, 3]. T2V-Zero [4] changes self-attention mechanisms to cross-frame attention. Tune-A-Video customizes video editing by adjusting attention weights [5]. Fate-Zero [6] keeps the video content intact via information from the inversion process. Animate-Zero [7] and VideoBooth [8] add layers that stress time.
@@
Figure 1: Existing methods suffer from (a) lack of fine-grained control, (b) imbalance between content fidelity and style richness, and (c) temporal inconsistency, whereas our UniVST effectively overcomes these challenges. The last row visualizes the coherence across frames using MimicMotion’s Y-T slices [9] for both BIVDiff [10] and our UniVST.
@@
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses the image or text to direct the style. StyleCrafter [11] personalizes video style with an extra training adapter; VideoBooth [8] uses the image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from a text prompt. The second method includes general video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, which is good for thematic expression but sometimes lacking in precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables varied scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
@@
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in blurring and alter the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As in BIVDiff [10] and AnyV2V [13], simply extending image stylization techniques to video often results in frame inconsistencies, causing various flickering and artifacts.
@@
In this paper, we propose UniVST, a unified localized [^1] Here, “localized” indicates a particular part of the video. video style transfer framework that addresses the above issues in a training-free manner. The first step in achieving this is to obtain masks of the region specified by users, which allows for precise control during the subsequent stylization process. However, unlike in the image domain [14, 15], where masks can be readily obtained using additional models, the video domain incurs additional inference overhead and presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. Recognizing the high redundancy between video frames, we propose a Point-Matching Mask Propagation strategy to efficiently extend the initial frame’s mask to subsequent frames. Leveraging feature maps from the inversion process [16], this strategy captures correlations in the latent space, eliminating the need for additional models.
By applying the inversion on both the original video and style image, we obtain the initial noise for each. Integrated with our point-matching mask propagation strategy, we can also obtain per-frame masks for the object region specified by users. Then, we establish new edited latents and construct a three-branch denoising architecture with the content, edited, and style branches, for training-free AdaIN-Guided Localized Video Stylization. It first utilizes localized latent blending to fuse information from the content branch and edited branch, achieving localized style control during the style transfer process. In addition, to enhance information interaction between edited and style branches during denoising, we propose a training-free AdaIN-guided latent-shift and attention-shift mechanism. Unlike the image stylization method [17], It effectively mitigates detail loss while adeptly balancing content fidelity and style richness during the stylization.
However, directly applying AdaIN-guided localized video stylization per-frame will introduce flickering and artifacts in final edited video. We empirically observe that this issue arises from a training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Inspired by previous works [2, 1], we further employ optical flow to enhance temporal consistency in the latent space. Specifically, we present a specialized Sliding-Window Consistent Smoothing scheme during denoising, which employs optical flow for spatial alignment at the pixel level and refines predicted noise in the latent space. As illustrated in Fig. 1, our smoothing scheme reduces flickering and artifacts while enhancing the temporal consistency in the final edited video.
Overall, our major contributions in this paper are as follows:
@@
  • We introduce, to the best of our knowledge, the first localized video stylization framework with diffusion models, featuring a novel point-matching mask propagation strategy.
@@
  • We develop a training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels. It effectively harmonizes content fidelity and style richness during the style transfer process.
@@
  • We present a sliding-window consistent smoothing scheme based on optical flow, which adeptly upholds temporal consistency throughout the style transfer process.
@@
  • Comprehensive experiments demonstrate that our framework outperforms several state-of-the-art methods across various qualitative and quantitative evaluation metrics.
@@ 2 Related Work
Image Style Transfer. Image style transfer aims to endow an image with a specific style, with early research focusing on non-diffusion methods. RSCT [18] calls for the separate transfer of foreground and background styles, combining them to ensure consistent styling. CAST [19] uses contrastive learning to extract style representations from image features, enhancing performance in arbitrary style transfer tasks. EFDM [20] utilizes exact feature distribution matching to align complex feature distributions across diverse style domains. With the development of diffusion models, recent studies have increasingly explored their application in style transfer. DiffStyle [21] employs LoRA [22] to refine image prompts for stylization, while ArtBank [23] enhances text embeddings through fine-tuning image prompts. StyleID [17] leverages key-value replacement coupled with an initial latent AdaIN [24] to preserve content fidelity. Z* [25] integrates content with diverse styles through an attention-based rearrangement mechanism. InstanceStylePlus [26] utilizes adapters to safeguard content retention and enhance stylistic expression. Although these methods demonstrate superior performance, directly applying them to video tasks using frame-wise processing often fails to ensure semantic continuity.
Video Style Transfer. Video style transfer prioritizes per-frame image style transfer while ensuring visual coherence. Similar to image style transfer, early research also concentrated on non-diffusion approaches. RT-NVST [27] uses feedforward networks for real-time video style transfer with a hybrid loss ensuring temporal consistency. RT-LVST [28] proposes a deep neural network for transferring artistic styles to specific video regions while preserving photorealism. These traditional non-diffusion models incur extra training costs and underperform diffusion models. Recent studies have increasingly employed diffusion models to accomplish video stylization. OCD [29] integrates object-centric sampling and merging, expediting the editing process. StyleCrafter [11] refines adapters using image prompts based on the T2V model. Style-A-Video [30] enhances stylistic alignment through strategic style guidance. Owing to the substantial computational demands of training, some studies have turned to training-free video stylization. Diffutoon [12] pioneers primary and editing branches for multi-segment editing in video colorization. BIVDiff [10] combines frame-by-frame image editing with temporal consistency modeling. AnyV2V [13] follows a similar initial editing strategy, subsequently leveraging an I2V model [31] to propagate changes across the video sequence. Although these diffusion-based methods exhibit strong performance, they struggle with fine-grained local editing while balancing content fidelity and style richness. In contrast, our proposed UniVST achieves fine-grained transfer while maintaining both content fidelity and style richness in a training-free manner.
Figure 2: Overall framework of our UniVST. It consists of three core components. (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Localized Video Stylization (Localized Latent Blending, AdaIN-Guided Latent-Shift and Attention-Shift) and (3) Sliding-Window Consistent Smoothing.
@@ 3 Preliminary
Latent Diffusion Model. LDM [32] adds and removes noise in a low-dim space by using the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a Gaussian noise $Z_{T}$, DDIM denoising [3] can be formulated as:
@@
in which ${Z}_{t\rightarrow 0}$ is an approximate estimation of $Z_{0}$ at time step $t$:
@@
where $\alpha_{t}$ is predefined parameter and $C=\phi$ for editing integrity.
@@
where $A_{t}$ and $B_{t}$ are the functions of time step $t$. During editing, the initial noise $Z_{T}$, which closely aligns with the original video distribution $Z_{0}$, can be accurately derived using this equation.
3D UNet Extension. Following previous works [34, 1], standard LDM will be extended into a 3D U-Net for video tasks. The $3\times 3$ convolutional kernels in convolutional blocks are expanded to $1\times 3\times 3$. To enable frame parallelization, the intermediate features are reshaped from $\mathbb{R}^{b\times c\times f\times h\times w}$ to $\mathbb{R}^{(bf)\times c\times h\times w}$ before self-attention and cross-attention. To enhance frame interaction, we follow T2V-Zero [4] to modify the self-attention into cross-frame attention:
$\begin{array}[]{c}Q^{i}=W^{query}\cdot\mathcal{F}^{i},\quad K^{i}=W^{key}\cdot[\mathcal{F}^{{1}},\mathcal{F}^{{i-1}}],\\ V^{i}=W^{value}\cdot[\mathcal{F}^{{1}},\mathcal{F}^{{i-1}}],\end{array}$ (4)
where $\mathcal{F}^{i}$ denotes the extracted feature of the $i$-th frame, while $W^{query}$, $W^{key}$, and $W^{value}$ represent the corresponding learnable mapping matrices, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicating concatenation operation.
4 Localized Video Style Transfer
Given an original video with $N$ frames $\{I^{i}\}_{i=1}^{N}$ and a style image $I^{s}$, our objective is to transform $\{I^{i}\}_{i=1}^{N}$ into a new sequence $\{\mathcal{I}^{i}\}_{i=1}^{N}$, which incorporates the reference style while preserving the primary objects’ original style. To achieve this, we propose UniVST, a unified framework for training-free localized video style transfer. The overall framework is illustrated in Fig. 2. UniVST first applies DDIM inversion to the original video and style image to obtain their initial noise and integrates Point-Matching Mask Propagation to generate masks for the object regions. It then performs AdaIN-Guided Localized Video Stylization with a three-branch architecture for information interaction. Moreover, Sliding-Window Consistent Smoothing is incorporated into the denoising process, enhancing the temporal consistency in the latent space.
Figure 3: The comparison between our mask propagation strategy and the traditional SAM model in terms of memory usage and inference time.
Algorithm 1 Efficient Point-Matching Mask Propagation
1:Input: $\mathcal{M}_{1}$ The mask for the first frame.
2:Input: $F_{t_{0}}$ The feature map from last up-sampling block.
@@ 4 Localized Video Style Transfer
6:Output: $\{\mathcal{M}^{i}\}_{i=1}^{N}$ The masks for all frames.
7:Initialize: PreviousFeatures [$\leftarrow$[\,]← [ ] Initialize empty list for previous features.
8:Initialize: PreviousMasks [$\leftarrow$[\,]← [ ] Initialize empty list for previous masks.
9:Initialize: FirstFeature [$\leftarrow$[\,]← [ ] Initialize empty list for first feature.
@@
11:    $\text{ForeIndices}\leftarrow\text{Where}(M_{i-1}=1)$
12:    $\text{ForeIndices}\leftarrow\text{Where}(M_{i-1}=0)$
13:    $\text{ForeIndices}\leftarrow\operatorname{RandomSample}(\text{ForeIndices},r\cdot\frac{\lvert\text{ForeIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
14:    $\text{BackIndices}\leftarrow\operatorname{RandomSample}(\text{BackIndices},r\cdot\frac{\lvert\text{BackIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
15:    $\text{CurrentIndices}\leftarrow\operatorname{Concat}(\text{ForeIndices},\text{BackIndices})$
16:    $\text{CurrentFeature}\leftarrow F_{t_{0}}^{i}[\text{CurrentIndices}]$
@@
18:       FirstFeature.append(CurrentFeature) Store the first frame.
19:       $\mathcal{M}^{i}\leftarrow\mathcal{M}^{1}$
@@
22:          PreviousFeatures.pop(0) Remove the oldest frame feature.
23:          PreviousMasks.pop(0) Remove the oldest frame mask.
@@
26:       $\mathcal{M}^{i}\leftarrow\text{KNN}(f,F_{t_{0}}^{i},\operatorname{Concat}(\text{PreviousMasks},\mathcal{M}^{1}),k)$
27:       PreviousFeatures.append(CurrentFeature) Store the current frame.
28:       PreviousMasks.append($\mathcal{M}_{i}$) Store the current mask.
@@
31:Return: $\{\mathcal{M}^{i}\}_{i=1}^{N}$